# Cross-correlation between the curvature perturbations and magnetic fields in pure ultra slow roll inflation

### v1

@@ V.2.2 Case of $p=4$
${\mathcal{B}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{27\,H_{{}_{\mathrm{I}}}^{6}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right],$ (68b)
${\mathcal{B}}_{3}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{16\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{3}\,k_{3}^{3}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left(k_{1}^{2}-k_{2}^{2}-k_{3}^{2}\right),$ (68c)
${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{243\,H_{{}_{\mathrm{I}}}^{6}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[k_{1}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+k_{2}^{4}+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right],$ (68d)
${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{27\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{3}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (68e)
${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=-\frac{9\,H_{{}_{\mathrm{I}}}^{7}}{16\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{5}\,(k_{1}+k_{2}+k_{3})^{5}\,\eta_{\mathrm{e}}}\,\left[2\,k_{3}^{2}\,(k_{1}^{2}+k_{2}^{2})+(k_{1}^{2}-k_{2}^{2})^{2}-3\,k_{3}^{4}\right]$
×$\quad\times\Biggl{\{}4\,k_{3}^{2}\,(7\,k_{1}^{2}+20\,k_{1}\,k_{2}+7\,k_{2}^{2})+\left[15\,k_{3}\,(k_{1}+k_{2})+3\,(k_{1}+k_{2})^{2}\right]\,(k_{1}^{2}+3\,k_{1}\,k_{2}+k_{2}^{2})$
$\quad+20\,k_{3}^{3}\,(k_{1}+k_{2})+4\,k_{3}^{4}\Biggr{\}}+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3}.$ (68f)
Note that, as in the case of $p=6$, it is the second and the fourth terms, viz. ${\mathcal{B}}_{2}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and ${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, that dominate the contributions. If we take into account these contributions, we find that the corresponding non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ turns out to be
$b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{15\,k_{1}}{16\,k_{2}^{2}\,k_{3}^{2}\,(k_{2}^{3}+k_{3}^{3})\,\eta_{\mathrm{e}}^{2}}\,\left[k_{1}^{4}+k_{2}^{4}+k_{3}^{4}-2\,k_{1}^{2}\,\left(k_{2}^{2}+k_{3}^{2}\right)+6\,k_{2}^{2}\,k_{3}^{2}\right].$ (69)
A notable feature of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in this case is the dependence on $\eta_{\mathrm{e}}$ as $\eta_{\mathrm{e}}^{-2}$. This is a large factor and it affects the amplitude of the parameter across all configurations. On inspecting the various limits, we find that, in the equilateral limit (i.e. when $k_{1}=k_{2}=k_{3}=k$), the expression (69) above simplifies to
@@ V.3 Consistency relation
Table 1: We have listed the values of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ obtained in the squeezed limit in the various scenarios of interest. We have assumed that $J\propto\eta^{-\alpha}$ in arriving at these results. We have also listed the corresponding values of $b_{{}_{\mathrm{NL}}}$ that we expect to obtain from the consistency condition (73). It is clear that, while the consistency relation is satisfied in slow roll inflation, it is violated in the ultra slow roll scenarios.
We have also listed the values of $b_{{}_{\mathrm{NL}}}^{\mathrm{cr}}$ that we are expected to obtain. From the table, we can clearly see that, while the consistency relation is satisfied in slow roll inflation, for a similar behavior of $J(\eta)$, it is violated in ultra slow roll inflation. The standard consistency relations arise in situations wherein the amplitude of the long wave length mode freezes [94, 95, 111, 112]. In slow roll inflation, we find that the consistency relation (73) is satisfied by the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the squeezed limit since, in such situations, the amplitude of the curvature perturbation goes to a constant value on super-Hubble scales. In contrast, in pure ultra slow roll inflation, the amplitude of the curvature perturbations grow indefinitely at late times. The violation of the consistency relation (73) that we encounter in the ultra slow roll scenarios can be ascribed to the growth of the curvature perturbations on super-Hubble scales. One expects the consistency relation (73) to be modified in such situations (for related discussions in the case of the scalar bispectrum, see Refs. [58, 57, 59]). It will be interesting to derive a generic consistency condition that governs the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the squeezed limit, including non-trivial evolution of the long wave length curvature perturbation. However, such a derivation is beyond the scope of this work and we hope to address the problem in the future.
@@ VI Conclusions
Models of inflation that permit a brief phase of ultra slow roll have attracted attention in the recent literature because they lead to scalar power spectra with higher amplitudes over small scales. Such scalar spectra with enhanced power on small scales can produce copious amounts of primordial black holes and also induce secondary gravitational waves of considerable strengths. In this work, we investigated the generation of magnetic fields in scenarios of ultra slow roll inflation. It has been pointed out that there arise certain challenges in generating magnetic fields with nearly scale invariant spectra and of observed strengths in single field inflationary models that permit a short period of ultra slow roll (see Ref. [53]; in this context, also see Ref. [54]). To circumvent such a challenge, rather than considering the electromagnetic action involving the popular non-conformal coupling function $J(\phi)$ [cf. Eq. (16)], we worked with an action involving a function of the form $J(X)$ [cf. Eq. (34)]. We focused on pure ultra slow roll scenarios. In other words, we assumed that the phase of ultra slow roll lasts throughout the entire duration of inflation. While such scenarios do not seem viable when compared with the observations, we considered such scenarios to gain a better understanding of the effects of ultra slow roll on the two-point and three-point functions involving the magnetic fields. We worked with a non-conformal coupling function $J(X)$ which turned out to be proportional to a power of the first slow roll parameter in the ultra slow roll scenarios. We find that, for suitable choices of parameters describing the non-conformal coupling function and the inflationary scenario, we can obtain scale invariant spectra for the magnetic fields.
@@
For our choice of $J(X)$, we also calculated the three-point cross-correlation ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ involving the curvature perturbation and the magnetic fields in the ultra slow roll scenarios. We examined the amplitude and shape of the associated non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ arising in a few different cases. When compared to slow roll inflation, we find that the values of the non-Gaussianity parameter in the ultra slow roll scenario of $p=6$, with $n=-2/3$, leading to scale invariant scalar and magnetic power spectra, are considerably smaller [cf. Eqs. (65), (66) and (67)]. However, we find that the values of $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ can vary significantly depending on the ultra slow roll scenario and the value of $n$ being considered. For example, in the case wherein we obtain a scale invariant spectrum for the magnetic field and a blue scalar power spectrum [with $\mathcal{P}_{{}_{\mathrm{S}}}(k)\propto k^{2}$], we obtain very large values for the non-Gaussianity parameter (see right panel of Fig. 1 and the accompanying discussion). In contrast, in the ultra slow roll scenario which leads to a scale invariant scalar power spectrum and a blue magnetic power spectrum [with $\mathcal{P}_{{}_{\mathrm{B}}}(k)\propto k^{2}$], we obtain very small values for the non-Gaussianity parameter (see App. D). Specifically, we had examined the validity of the consistency condition that the three-point function of interest is expected to satisfy in the squeezed limit [cf. Eq. (73)]. This consistency relation is known to be valid in the the standard slow roll inflationary scenarios (see App. D for a brief discussion on this point). But, we observe that, in all the ultra slow roll scenarios that we have considered, the aforementioned consistency relation is violated. This can be attributed to the fact that, in the ultra slow roll scenarios, in contrast to slow roll inflation, the curvature perturbation grows strongly on super-Hubble scales.
@@
It would be interesting to investigate the manner in which the super-Hubble evolution of the long wavelength mode (the curvature perturbation in this case) modifies the consistency relation (for discussions on the scalar bispectrum in such situations, see, for instance, Refs. [58, 59]). Rather than the pure ultra slow roll models, it would be compelling to study the production of magnetic fields in the inflationary models involving a brief phase of ultra slow roll sandwiched by two epochs of slow roll inflation. Such scenarios are in better agreement with the observations and would hence provide us with testable predictions. We defer these interesting possibilities to a future work.
@@ Acknowledgements.
ST thanks the Indian Institute of Astrophysics, Bengaluru, India, for financial support through a postdoctoral fellowship. DC would like to thank the Indian Institute of Astrophysics, Bengaluru, India, and the Department of Science and Technology, Government of India, for support through the INSPIRE Faculty Fellowship grant DST/INSPIRE/04/2023/000110. DC would also like to thank the Indian Institute of Science, Bengaluru, India, for support through the C. V. Raman postdoctoral fellowship. HVR thanks the Raman Research Institute, Bengaluru, India, for support through a postdoctoral fellowship. LS wishes to thank the Indo-French Centre for the Promotion of Advanced Research for support of the proposal 6704-4 under the Collaborative Scientific Research Programme.

### v2

@@ V.2.2 Case of $p=4$
${\mathcal{B}}_{2}(\bm{k}_{1},\bm{k}_{2},\bm{k}_{3})$ $=\frac{9\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{16\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{3}\,k_{3}^{3}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left(k_{1}^{2}-k_{2}^{2}-k_{3}^{2}\right),$ (68b)
${\mathcal{B}}_{3}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{27\,H_{{}_{\mathrm{I}}}^{6}\,\eta_{\mathrm{i}}^{2}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{3}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{2}\,\left[(k_{1}^{2}-k_{2}^{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (68c)
${\mathcal{B}}_{4}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{27\,\pi\,H_{{}_{\mathrm{I}}}^{6}}{64\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{1}^{3}\,k_{2}^{3}\,k_{3}^{5}\,\eta_{\mathrm{e}}}\left[(k_{1}^{2}-k_{2}^{2})^{2}+2\,(k_{1}^{2}+k_{2}^{2})\,k_{3}^{2}-3\,k_{3}^{4}\right]+{\bm{k}}_{2}\leftrightarrow{\bm{k}}_{3},$ (68d)
${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{27\,H_{{}_{\mathrm{I}}}^{6}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right],$ (68e)
${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{243\,H_{{}_{\mathrm{I}}}^{6}}{128\,M_{{}_{\mathrm{Pl}}}^{2}\,k_{2}^{5}\,k_{3}^{5}\,\epsilon_{1\mathrm{i}}}\,\left(\frac{\eta_{\mathrm{i}}}{\eta_{\mathrm{e}}}\right)^{4}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right].$ (68f)
Note that, as in the case of $p=6$, it is the fifth and the sixth terms, viz. ${\mathcal{B}}_{5}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ and ${\mathcal{B}}_{6}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$, that dominate the contributions. If we take into account these contributions, we find that the corresponding non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ turns out to be
$b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ $=\frac{15\,k_{1}}{16\,k_{2}^{2}\,k_{3}^{2}\,(k_{2}^{3}+k_{3}^{3})\,\eta_{\mathrm{e}}^{2}}\,\left[k_{1}^{4}+k_{2}^{4}-2\,k_{1}^{2}\,(k_{2}^{2}+k_{3}^{2})+6\,k_{2}^{2}\,k_{3}^{2}+k_{3}^{4}\right].$ (69)
A notable feature of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in this case is the dependence on $\eta_{\mathrm{e}}$ as $\eta_{\mathrm{e}}^{-2}$. This is a large factor and it affects the amplitude of the parameter across all configurations. On inspecting the various limits, we find that, in the equilateral limit (i.e. when $k_{1}=k_{2}=k_{3}=k$), the expression Eq. (69) above simplifies to
@@ V.3 Consistency relation
Table 1: We have listed the values of the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ obtained in the squeezed limit in the various scenarios of interest. We have assumed that $J\propto\eta^{-\alpha}$ in arriving at these results. We have also listed the corresponding values of $b_{{}_{\mathrm{NL}}}$ that we expect to obtain from the consistency condition given in Eq. (73). It is clear that, while the consistency relation is satisfied in slow roll inflation, it is violated in the ultra slow roll scenarios.
We have also listed the values of $b_{{}_{\mathrm{NL}}}^{\mathrm{cr}}$ that we are expected to obtain. From the table, we can clearly see that, while the consistency relation is satisfied in slow roll inflation, for a similar behavior of $J(\eta)$, it is violated in ultra slow roll inflation. The standard consistency relations arise in situations wherein the amplitude of the long wave length mode freezes [96, 97, 113, 114]. In slow roll inflation, we find that the consistency relation in Eq. (73) is satisfied by the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the squeezed limit since, in such situations, the amplitude of the curvature perturbation goes to a constant value on super-Hubble scales. In contrast, in pure ultra slow roll inflation, the amplitude of the curvature perturbations grows indefinitely at late times. The violation of the consistency relation [as given by Eq. (73)] that we encounter in the ultra slow roll scenarios can be ascribed to the growth in the amplitude of the curvature perturbations on super-Hubble scales. One expects the consistency relation in Eq. (73) to be modified in such situations (for related discussions in the case of the scalar bispectrum, see Refs. [60, 59, 61]). It will be interesting to derive a generic consistency condition that governs the non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ in the squeezed limit, including non-trivial evolution of the long wave length curvature perturbation. However, such a derivation is beyond the scope of this work and we hope to address the problem in the future.
@@ VI Conclusions
Models of inflation that permit a brief phase of ultra slow roll have attracted attention in the recent literature because they lead to scalar power spectra with higher amplitudes over small scales. Such scalar spectra with enhanced power on small scales can produce copious amounts of primordial black holes and also induce secondary gravitational waves of considerable strengths. In this work, we investigated the generation of magnetic fields in scenarios of ultra slow roll inflation. It has been pointed out that there arise certain challenges in generating magnetic fields with nearly scale invariant spectra and of observed strengths in single field inflationary models that permit a short period of ultra slow roll (see Ref. [55]; in this context, also see Ref. [56]). To circumvent such a challenge, rather than considering the electromagnetic action involving the popular non-conformal coupling function $J(\phi)$ [cf. Eq. (16)], we worked with an action involving a function of the form $J(X)$ [cf. Eq. (34)]. We focused on pure ultra slow roll scenarios. In other words, we assumed that the phase of ultra slow roll lasts throughout the entire duration of inflation. While such scenarios do not seem viable when compared with the observations, we considered these scenarios to gain a better understanding of the effects of ultra slow roll on the two-point and three-point functions involving the magnetic fields. We worked with a non-conformal coupling function $J(X)$ which turned out to be proportional to a power of the first slow roll parameter in the ultra slow roll scenarios. We find that, for suitable choices of parameters describing the non-conformal coupling function and the inflationary scenario, we can obtain scale invariant spectra for the magnetic fields. However, we should mention that, for our choice of parameters, the coupling function does indeed exhibit the strong coupling problem during early times. One of the ways to ameliorate this issue is to consider a ‘sawtooth’ coupling (for detailed discussions, see, for example, Refs. [115, 103, 116]). We are currently studying implementations of such coupling functions in ultra slow roll scenarios.
@@
For our choice of $J(X)$, we also calculated the three-point cross-correlation ${\mathcal{B}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ involving the curvature perturbation and the magnetic fields in the ultra slow roll scenarios. We examined the amplitude and shape of the associated non-Gaussianity parameter $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ arising in a few different cases. When compared to slow roll inflation, we find that the values of the non-Gaussianity parameter in the ultra slow roll scenario of $p=6$, with $n=-2/3$, leading to scale invariant scalar and magnetic power spectra, are considerably smaller [cf. Eqs. (65), (66) and (67)]. However, we find that the values of $b_{{}_{\mathrm{NL}}}({\bm{k}}_{1},{\bm{k}}_{2},{\bm{k}}_{3})$ can vary significantly depending on the ultra slow roll scenario and the value of $n$ being considered. For example, in the case wherein we obtain a scale invariant spectrum for the magnetic field and a blue scalar power spectrum [with $\mathcal{P}_{{}_{\mathrm{S}}}(k)\propto k^{2}$], we obtain very large values for the non-Gaussianity parameter (see right panel of Fig. 1 and the accompanying discussion). In contrast, in the ultra slow roll scenario which leads to a scale invariant scalar power spectrum and a blue magnetic power spectrum [with $\mathcal{P}_{{}_{\mathrm{B}}}(k)\propto k^{2}$], we obtain very small values for the non-Gaussianity parameter (see App. D). Specifically, we had examined the validity of the consistency condition that the three-point function of interest is expected to satisfy in the squeezed limit [cf. Eq. (73)]. This consistency relation is known to be valid in the standard slow roll inflationary scenarios (see App. D for a brief discussion on this point). But, we observe that, in all the ultra slow roll scenarios that we have considered, the aforementioned consistency relation is violated. This can be attributed to the fact that, in the ultra slow roll scenarios, in contrast to slow roll inflation, the amplitude of the curvature perturbation grows strongly on super-Hubble scales.
@@
It would be interesting to investigate the manner in which the super-Hubble evolution of the long wavelength mode (the curvature perturbation in this case) modifies the consistency relation (for discussions on the scalar bispectrum in such situations, see, for instance, Refs. [60, 61]). Rather than the pure ultra slow roll scenarios, it would be compelling to study the production of magnetic fields in inflationary models involving a brief phase of ultra slow roll sandwiched by two epochs of slow roll inflation. Such scenarios are in better agreement with the observations and would hence provide us with testable predictions. We defer these interesting possibilities to a future work.
@@ Acknowledgements.
ST thanks the Indian Institute of Astrophysics, Bengaluru, India, for financial support through a postdoctoral fellowship. DC would like to thank the Indian Institute of Astrophysics, Bengaluru, India, and the Department of Science and Technology, Government of India, for support through the INSPIRE Faculty Fellowship grant DST/INSPIRE/04/2023/000110. DC would also like to thank the Indian Institute of Science, Bengaluru, India, for support through the C. V. Raman postdoctoral fellowship. HVR thanks the Raman Research Institute, Bengaluru, India, for support through a postdoctoral fellowship. HVR also acknowledges support by the MUR PRIN2022 Project “BROWSEPOL: Beyond standaRd mOdel With coSmic microwavE background POLarization”-2022EJNZ53 financed by the European Union—Next Generation EU. LS wishes to thank the Indo-French Centre for the Promotion of Advanced Research for support of the proposal 6704-4 under the Collaborative Scientific Research Programme.